OpenStudy (anonymous):

A wire is 1 foot long and is cut into two pieces. One piece is bent into a square and the other is bent into a semicircle. Let x= the amount used for the square and let A(x) represent the sum of the total enclosed areas of square and semicircle. Find the maximum and minimum values of A(x).

OpenStudy (anonymous):

a negative area cannot exist however that is a verified derivative. did i go wrong or is it true that a max area would occur if he used the wire for only a semicircle?

OpenStudy (anonymous):

my circle area is wrong...working on it

OpenStudy (anonymous):

still obtaining a negative area ... for d/dx

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

I have not gone very far, but the area you gave for the semi-circle is actually that for the whole circle. Can you check if that was the only problem?

OpenStudy (anonymous):

updated

OpenStudy (anonymous):

A(8PI/(2PI+4)=-0.0678322

OpenStudy (mathmate):

let me see, you're good up to r=...

OpenStudy (mathmate):

But since the area of the semi-circle is \(\pi r^2/2\), I think you forgot to square the radius.

OpenStudy (anonymous):

fixed the negative area. but a semicircle still provides more area.

OpenStudy (mathmate):

The A'(x) is still not up-to-date! :(

OpenStudy (mathmate):

I got: \(\huge A'(x)=\frac{x}{8}-\frac{\pi(1-x)}{(\pi+2)^2}\) Hope that helps!

OpenStudy (anonymous):

how did the square stay at the bottom?

OpenStudy (mathmate):

The square stays at the bottom because it is a constant, so taking derivative does not affect it.

OpenStudy (anonymous):

this?

OpenStudy (mathmate):

Almost! You forgot there is an x on the right hand side when you isolated x.

OpenStudy (mathmate):

I would cross multiply and solve for x. \((\pi+2)^2x = 8\pi (1-x)\) expand : \((\pi+2)^2x = 8\pi -8\pi x)\) And \(x = \frac{8\pi}{(\pi+2)^2+8\pi}\) => \(\large x = \frac{8\pi}{\pi^2+12\pi+4}\)

OpenStudy (anonymous):

better answer but still holds true that a semicircle maxmizes area

OpenStudy (mathmate):

Yes, you need both max and min.

OpenStudy (anonymous):

so the derivative is the min :)

OpenStudy (mathmate):

You can find out if it is a max or min by finding the second derivative. If the second derivative is >0, then it is a minimum, and if it is negative, it is a maximum. Here the second derivative is \(\frac{\pi}{(\pi+2)^2}+\frac{1}{8}=0.24>0\), so it is a minimum. The other two values you need to check are the boundaries, i.e. x=0, and x=1.

OpenStudy (mathmate):

Did you calculate the area corresponding to the minimum?

OpenStudy (anonymous):

these are tough and very involved. especially in summer school with a very very bad professor. im beginning to lose steam on this. how am i so far

OpenStudy (anonymous):

A(0) being max and A(.24) min

OpenStudy (mathmate):

I have xmin=8pi/.... = 0.487... and A(0.487...)=0.03046... A(0)=0.0594... (semi-circle) A(1)=0.0625 (square)

OpenStudy (anonymous):

thank you very much for your help. i really appreciate the time youve given me. it appears though that i need to do more of these. i find it very frustrating understanding the concepts but constantly suffering from these small errors that add up and are costly and detrimental to my grade. ive spent 20-30 hrs a week on this stuff and i barely have a 70%. with only two weeks left in then summer course. with a 30 hr work week. i wish my professor didnt suck

OpenStudy (mathmate):

Yes, I agree! Like everything else, practice makes perfect. Summer courses are stressful, because there is less elapsed time to digest the information, and of course assignments are on your back. It is good that you spend an appreciable amount of time to work on the material. Keep working at it and you'll get your reward eventually. One way to avoid the nuisance of costly small errors is to review your work at every step. While you're fresh at it, it takes very little time. If you have to find a mistake at the end, it will seem a monumental task. Best of luck to your summer course, and of course, post any time! mathmate